You plan to put a fence around a rectangular lot. The length of the lot must be at least 70ft. The cost for the fence along the length of the lot is $2 per foot, and the cost of the fence along the width is $3 per foot. The total cost can not exceed $360.
a. Use two variables to write a system of inequalities that models the problem
b. Graph the system, and shade the feasible region.
c.What is the maximum width of the lot if the length is 60 feet?

Answer by AnlytcPhil(1806)

You plan to put a fence around a rectangular lot. The length of the lot must be
at least 70ft. The cost for the fence along the length of the lot is $2 per
foot, and the cost of the fence along the width is $3 per foot. The total cost
can not exceed $360. 
a. Use two variables to write a system of inequalities that models the problem

Let x = the length and y = the width
      
 ___________
|     x     | 
|           |y
|           |
|___________|

>>...The length of the lot must be at least 70ft...<<

Translation:  x >= 70

>>...The cost for the fence along the length of the lot is $2 per foot,...<<

Translation: cost along the two lengths is $2 times 2x (there are two
lengths). This amounts to 2(2x) or 4x.

>>...the cost of the fence along the width is $3 per foot...<<

Translation: cost along the two widths is $3 times 2y (there are two
widths). This amounts to 3(2y) or 6y.

>>...The total cost can not exceed $360...<<

Translation 4x + 6y <= 360

We must also require that the variables x and y are not negative. x >= 70
meets this requirement, so we only need y >= 0

x >= 70
4x + 6y <= 360 
y >= 0

These are the inequalities which model the problem.

---------------------------------------------------

b. Graph the system, and shade the feasible region.

Draw the graphs of the boundary lines

x = 70           a vertical line through (70,0)
4x + 6y = 360    a slanted line from (0,60) to (90,0)
y = 0            the x-axis

The feasible region is a triangle whose corners are

(70, 0), (70, 13 1/3) and (90,0)

Shade this triangle. 

----------------------------------------------------------------

c.What is the maximum width of the lot if the length is 60 feet?

y is the width and the maximum value must occur at one of the corner
points.  Thus the maximum value of y is at the corner point (70, 13 1/3),
so the maximum width is 13 1/3 feet.  (A very narrow lot, really a strip)

Edwin
AnlytcPhil@aol.com